Show that the function f defined on [1,0] by f(x) = (-1)n-1 for 1/n+1 <x< = 1/n (for n= 1,2,3,...) is integrable on [0,1]


1
Expert's answer
2022-06-01T13:34:43-0400

ANSWER

Since the set of value of the function is "\left \{ -1,1 \right \}" . then the function is bounded

"|f(x)| = 1" for all "x\in [-1,1]" .

In each interval "(\frac {1}{n+1}, \frac {1}{n})" "(n\in \N)" the function is constant "(=-1" or "1)" , therefore it is continuous. At the points "x_{n}=\frac{1}{n}" the function is discontinuous , because

"\lim_{x\rightarrow (\frac {1}{2n-1})^{-} }f(x) = \lim_{x\rightarrow (\frac{1}{2n-1})^{-} } (-1)^{ (2n-2) }=1 \\\lim_{x\rightarrow (\frac {1}{2n-1})^{+} }f(x) = \lim_{x\rightarrow (\frac{1}{2n-1})^{+} } (-1)^{ (2n-3) }=-1"

and vice versa for points "x_{2n}" . Note, that "\lim_{n\rightarrow \infty }x_{n}=0," and "\lim_{n\rightarrow \infty}f(x_{n})=\lim_{n\rightarrow \infty}(-1)^{n}" does not exist. Therefore "x=0" is also a point of discontinuity of function.

Denote "B" the set of discontinuity of function. "B= \left \{ 0,1,\frac{1}{2} ,..., \frac{1}{n},...\right \}" and use

Definition .A set "A\subset \R" has measure (Lebesgue) zero "(m(A)=0)" , if for all "\epsilon>0," there is a countable collection of open intervals "\left \{ I_{0} .I_{1},..., I_{n},...\right \}" such that

"A\subset\bigcup_{n=0}^{\infty} I_{n}" and . "\sum_{n=0}^{\infty} m(I_{n} )<\epsilon" .

If "I=(\alpha,\beta)" , then "m(I)=\beta-\alpha."

Let "A=B, I_{0}=(-\frac{\epsilon}{4}, \frac{\epsilon}{4}), I_{n}=(\frac{1}{n}-\frac{\epsilon}{2^{n+2}}, \frac{1}{n}+\frac{\epsilon}{2^{n+2}})" . "0\in I_{0}, \frac{1}{n}\in I_{n}." So "B\subset\bigcup_{n=0}^{\infty} I_{n}" ,

"m(I_{n})=\frac{\epsilon}{2^{n+1}} (n=0,1,2,... )" and "\sum_{n=0}^{\infty} m(I_{n} )=\sum_{n=0}^{\infty} \frac{\epsilon}{2^{n+1}} =\frac{\frac{\epsilon}{2}}{1-\frac{1}{2}} =\epsilon" . Hence , "m(B)=0."

Therefore, by the Riemann -Lebesgue Theorem, "f" is Riemann integrable on "[0,1]".